OpenStudy (anonymous):

Urgent Help Please : A ladder leans against a building. The angle of elevation of the ladder is 70°. The top of the ladder is 25 ft from the ground. To the nearest tenth of a foot, how far from the building is the base of the ladder? 20.5 ft 30.5 ft 32.3 ft 39.5 ft

OpenStudy (anonymous):

@IrishBoy123

OpenStudy (anonymous):

tan70=25/x 2.74=25/x x=9.09 ft

OpenStudy (anonymous):

It is not an answer choice though

OpenStudy (irishboy123):

draw it first @utcung2103's approach made a lot of sense